# Quantum anomalous Hall phases in gated rhombohedral graphene

### v1

@@ Appendix A Appendix
Consider an interface Hamiltonian $H_{I}$ with matrix-valued symbol $a(\mathbf{x},{\mathbf{k}})$ in the Weyl quantization, i.e.,
@@
with $a(\mathbf{x},{\mathbf{k}})$ satisfying hypothesis [H1] in [26, 8] of order $m=1$, i.e., such that $a(\mathbf{x},{\mathbf{k}})$ has singular values bounded above and below by $O(|{\mathbf{k}}|)$ and such that $a(\mathbf{x},{\mathbf{k}})=a^{N}({\mathbf{k}})$ for $y\geq R>0$ while $a(\mathbf{x},{\mathbf{k}})=a^{S}({\mathbf{k}})$ for $y\leq-R$. We further assume that $a^{N/S}({\mathbf{k}})$ are gapped for all ${\mathbf{k}}\in\mathbb{R}^{2}$ in the energy interval $(-E_{0},E_{0})$, i.e., all of their eigenvalues lie outside of this interval. We use the notation $\mathbf{x}=(x,y)$ while ${\mathbf{k}}=(k_{x},k_{y})$.
Let $P=P(x)$ be a function that depends only the spatial coordinate $x$ with $P(x)=0$ for $x<x_{0}-\delta$ and $P(x)=1$ for $x>x_{0}+\delta$ for some $x_{0}\in\mathbb{R}$ and $\delta>0$. The function $P(x)$ should be interpreted as the observable quantifying the field density in the (right) half-space $x\geq x_{0}$. Let $0\leq\varphi\in C^{\infty}(\mathbb{R})$ be a function such that $\varphi(E)=0$ for $E\leq-E_{0}$ and $\varphi(E)=1$ for $E\geq E_{0}$. Thus $\varphi^{\prime}(H_{I})$ defines a density of states that cannot propagate into the N and S insulating (for such energies) bulks.
The operator $i[H,P]$ may be interpreted as a current operator of excitations crossing the vertical line $x=x_{0}$ per unit time. The expectation of this operator against the density $\varphi^{\prime}(H)$ is then defined as
$\sigma_{I}[H_{I}]={\rm Tr}\ i[H_{I},P]\varphi^{\prime}(H_{I}).$ (22)
This assumes that $i[H_{I},P]\varphi^{\prime}(H_{I})$ is a trace-class operator, which is indeed the case since Hypothesis [H1] holds [3, 8, 26]. This current is quantized: $2\pi\sigma_{I}[H_{I}]\in\mathbb{Z}$, reflecting the topological nature of the edge states and the robustness of this current to perturbations of $a(\mathbf{x},{\mathbf{k}})$ satisfying condition [H1] [3, 8, 26].
@@
where $\Pi^{h}_{j}({\mathbf{k}})=\psi^{h}_{j}({\mathbf{k}})\otimes\psi^{h}_{j}({\mathbf{k}})$ are rank-one projectors and $\lambda^{h}_{j}(\mathbf{k})$ are the corresponding eigenvalues. In the RHG application, all bands are simple and hence in fact real-analytic in $\mathbf{k}$}. Associated to each (arbitrary-rank) projector family $\mathbb{R}^{2}\ni{\mathbf{k}}\mapsto\Pi({\mathbf{k}})$ is the following integral of the associated (Berry) curvature
@@
Here $\rm tr$} stands for standard matrix trace. Because $\mathbb{R}^{2}$ is not a compact manifold, the above integral is not necessarily a (stable with respect to continuous deformations) Chern number or even necessarily integral-valued. In fact, in many applications, the above integral takes an arbitrary continuum of values (see [6] for a model of bilayer graphene). One reason for this fact is that the projector $\Pi$ may not have a constant value as $k\to\infty$. In such a setting, we may still be able glue two projectors, one from $H^{N}$ and the other one from $H^{S}$ by radial compactification of two Euclidean planes on the Riemann sphere [3, 12, 27], and obtain a well-defined Chern number. We recall the main steps of the procedure.
@@
That the gluing condition (25) holds for the RHG model (21) is a consequence of ellipticity: as $|{\mathbf{k}}|\to\infty$, the ranges of $P^{N}_{\ell}$ and $P^{S}_{\ell}$ are independent of the parameters $(\varepsilon,\gamma,u_{j})$ and equal. Moreover, after replacing each term $k\sigma_{1}$ by $\alpha_{n}k\sigma_{1}$ as was done in the proof of Theorem 2.3, it is then straightforward to observe that the eigenvectors obtained in the limit $k\to\infty$ are independent of the displacement fields $\pm u$ and hence the gluing conditions obtained for each band separately.
Bulk edge correspondence
@@
Acknowledgments
The authors acknowledge stimulating discussions with Allan MacDonald on graphene systems and Jeremy Hoskins for invaluable advice and input regarding numerical techniques. This work was funded in part by NSF grant DMS-230641.

### v2

@@ Appendix A Classification of elliptic Hamiltonians
Consider an interface Hamiltonian $H_{I}$ with matrix-valued symbol $a(\mathbf{x},{\mathbf{k}}):\mathbb{R}^{4}\to\mathbb{C}^{2m\times 2m}$ in the Weyl quantization, i.e.,
@@
with $a(\mathbf{x},{\mathbf{k}})$ satisfying hypothesis [H1] in [30, 6] of order $\langle\mathbf{x}\rangle$ (of order $m=1$ in the notation of [30, 6]), i.e. such that the minimum singular value of $a(\mathbf{x},{\mathbf{k}})$, $a_{min}(\mathbf{x},{\mathbf{k}})$ is bounded below according to $a_{min}(\mathbf{x},{\mathbf{k}})\geq C|{\mathbf{k}}|-1$ for some $C>0$ and such that $a(\mathbf{x},{\mathbf{k}})=a^{N}({\mathbf{k}})$ for $y\geq R>0$ while $a(\mathbf{x},{\mathbf{k}})=a^{S}({\mathbf{k}})$ for $y\leq-R$. As such $H_{I}$ is an unbounded operator whose domain must be $H^{1}(\mathbb{R}^{2};\mathbb{C}^{2m})$ or a subset thereof. Fixing $H^{1}(\mathbb{R}^{2};\mathbb{C}^{2m})$ as the domain of $H_{I}$ ensures self-adjointness in the main part of the paper. We further assume that $a^{N/S}({\mathbf{k}})$ are gapped for all ${\mathbf{k}}\in\mathbb{R}^{2}$ in the energy interval $(-E_{0},E_{0})$, i.e., all of their eigenvalues lie outside of this interval. We use the notation $\mathbf{x}=(x,y)$ while ${\mathbf{k}}=(k_{x},k_{y})$.
Let $F=F(x)\in C^{\infty}(\mathbb{R})$ be a function that depends only the spatial coordinate $x$ with $F(x)=0$ for $x<x_{0}-\delta$ and $F(x)=1$ for $x>x_{0}+\delta$ for some $x_{0}\in\mathbb{R}$ and $\delta>0$. The function $F(x)$ should be interpreted as the observable quantifying the field density in the (right) half-space $x\geq x_{0}$. Let $0\leq\varphi\in C^{\infty}(\mathbb{R})$ be a function such that $\varphi(E)=0$ for $E\leq-E_{0}$ and $\varphi(E)=1$ for $E\geq E_{0}$. Thus $\varphi^{\prime}(H_{I})$ defines a density of states that cannot propagate into the N and S insulating (for such energies) bulks.
The operator $i[H_{I},F]$ may be interpreted as a current operator of excitations crossing the vertical line $x=x_{0}$ per unit time. The expectation of this operator within the appropriate Hilbert space (in our case $H^{1}(\mathbb{R}^{2};\mathbb{C}^{2m})$) against the density $\varphi^{\prime}(H)$ is then defined as
$\sigma_{I}[H_{I}]={\rm Tr}\ i[H_{I},F]\varphi^{\prime}(H_{I}),$ (24)
Tr denoting the trace in the appropriate Hilbert space. This assumes that $i[H_{I},F]\varphi^{\prime}(H_{I})$ is a trace-class operator, which is indeed the case whenever [H1] holds [6, 4, 30]. Note however that if $\varphi^{\prime}$ is supported on any interval containing bulk spectrum, $\sigma_{I}[H_{I}]$ is clearly not trace-class so that $\sigma_{I}$ necessarily characterizes a bulk band gap. This current is quantized: $2\pi\sigma_{I}[H_{I}]\in\mathbb{Z}$, reflecting the topological nature of the edge states and the robustness of this current to perturbations of $a(\mathbf{x},{\mathbf{k}})$ satisfying condition [H1] [6, 4, 30].
@@
where $\Pi^{h}_{j}({\mathbf{k}})=|\psi^{h}_{j}({\mathbf{k}})\rangle\langle\psi^{h}_{j}({\mathbf{k}})|$ are rank-one projectors and $\lambda^{h}_{j}(\mathbf{k})$ are the corresponding eigenvalues. The spectral theorem ensures such a decomposition exists for self-adjoint operators. In the RHG application, all bands are simple and hence in fact real-analytic in $\mathbf{k}$}. Associated to each (arbitrary-rank) projector family $\mathbb{R}^{2}\ni{\mathbf{k}}\mapsto\Pi({\mathbf{k}})$ is the following integral of the associated (Berry) curvature
@@
Here $\rm tr$} stands for standard matrix trace. Because $\mathbb{R}^{2}$ is not a compact manifold, the above integral is not necessarily a (stable with respect to continuous deformations) Chern number or even necessarily integral-valued. In fact, in many applications, the above integral takes an arbitrary continuum of values (see [2] for a model of bilayer graphene). One reason for this fact is that the projector $\Pi$ may not have a constant value as $k\to\infty$. In such a setting, we may still be able glue two projectors, one from $H^{N}$ and the other one from $H^{S}$, by radial compactification of two Euclidean planes on the Riemann sphere [6, 16, 31], and obtain a well-defined Chern number. We recall the main steps of the procedure.
@@
That the gluing condition (27) holds for the RHG model (23) is a consequence of ellipticity: as $|{\mathbf{k}}|\to\infty$, the ranges of $P^{N}_{\ell}$ and $P^{S}_{\ell}$ are independent of the parameters $(\gamma,u)$ and equal. Moreover, after replacing each term $k\sigma_{1}$ by $\alpha_{n}k\sigma_{1}$ as was done in the proof of Theorem 2.3, it is then straightforward to observe that the eigenvectors obtained in the limit $k\to\infty$ are independent of the displacement fields $\pm u$ and hence the gluing conditions obtained for each band separately.
Bulk edge correspondence.